INAF queue observing, 2013 May 7-8

Observer: F. Cusano, R. Speziali
Telescope Operator: S. Allanson
Support Astronomer: O. Kuhn
Instrument: LBC

Summary

Partially cloudy at the sunset and strong wind gusts (in between 12 and 23 m/s). For the strong wind the dome remain closed up to 04:40 UT.
There are some problems with the secondary mirror so we decide to not use MODS in this night.
The sky is clear at the beginning of the observations (04:40 UT). Average seeing 1.1 arcsec during the observations.
At 07:24 UT dome closed for high humidity.
We open again the dome at 09:17, but the conditions are not optimal (seeing from 1.2 to 1.6 arcsec, some wind gusts).

Data Summary

Project_RUN Objects FilterSorted descending ExpTime Notes
Grazian_30A GOODS-North Uspec, r-sloan 1.5 h seeing ~ 1.1
" " U-bessel, V-bessel 50s x 40s  
Piotto_34B M13 U-bessel, i-sloan 50s x 40s  
Photometric STD SA107 U, B, Us, g; V, r, i, r 12m airmass 1.3
" " B-bessel, V-bessel 20s x 20s, 10.7 m x 9.7 m  
calibration sktflat B-bessel, Uspec; V-bessel,Y    

Details

00:40 Start 25Bias_BINO

03:00 The dome stays closed for the strong wind gusts

04:32 The TO opens the dome, the wind velocity is decreasing.

04:41 Executing the program Grazian_30A, launched grazian_pointing.ob

04:42 dofpia, /first, /X2

04:49 Seeing ~1.2 arcsec, launched again grazian_pointing.ob, lbcrangebal with 045027 and 045022

04:56 dofpia, /first, /X2

05:01 Launched grazian_GOODSN_UR_34.ob (seeing ~1.2 arcsec)
-elongated stars in some images of the blue channel
-stopped at 18/20 frames, in the blue channel the images are out of focus (the blue primary mirror has some temperature gradient)

05:41 dofpia, /X2

05:46 Launched grazian_GOODS_UR_35.ob (seeing 1.1-1.3)

06:26 dofpia, /X2

06:30 Launched grazian_GOODS_UR_36.ob (seeing 1, at 07:05 difference between dewpoint and air temperature smaller than 2 degrees C)

07:14 Slew to standard field SA107

07:17 dofpia, /X2

07:24 The TO has to close the dome, humidity 96%

09:17 Weather conditions are better the TO opens the dome

09:24 slew to standard field SA107

09:27 dofpia, /first, /X2

09:37 Launched SA107_UBUsgVrir.ob

09:50 slewing to scientific target of Brocato_32C (Draco)

09:51 dofpia, /X2

09:56 Launched Dra_pointing.ob, stopped; seeing to high 1.2-1.4 for Brocato_32C, changing target

09:59 Slewing to Piotto_34B (M13)

10:00 dofpia, /X2

10:05 Launched M13_Ui_1Bri.ob, ob finished but there was an error in the ob (the red channel was setted to filter V, while it was supposed to be on filter i), we corrected the ob and launched again (seeing >1.4)

10:22 dofpia

10:35 Launched M13_BV_1Bri.ob

10:40 Launched M13_BV_1.ob

10:48 Launched M13_BV_2.ob

11:04 Launched M13_BV_3.ob, stopped objects too much elongated

11:11 dofpia

11:21 Launched again M13_BV_3.ob (twilight)

11:35 starting skyflat (too early)

11:52 skyflat_BV_r1806d+00_pa0_5.ob scale exptimes 4

11:57 skyflat_BV_r1806d+00_pa0_5.ob scale exptimes 1.5 (stopped)

11:58 skyflat_SY_r1806d+00_pa0_5.ob scale exptimes 1.0

12:09 dome closed

-- FeliceCusano - 08 May 2013
Topic revision: r1 - 31 Jul 2024, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback